OpenStudy (anonymous):

Simplify the expression (3/4x -1/2) - (1/4x- 3/2)

OpenStudy (anonymous):

MULTIPLY OUT the minus sign. and you get: 3/4x - 1/2 - 1/4x + 3/2 then do simple addition/subtraction and you get: 2/4x + 2/2 then simplify, and you get: 1/2x + 1

OpenStudy (anonymous):

((3)/(4)*x-(1)/(2))-((1)/(4)*x-(3)/(2)) Multiply (3)/(4) by x to get (3x)/(4). ((3x)/(4)-(1)/(2))-((1)/(4)*x-(3)/(2)) Multiply (1)/(4) by x to get (x)/(4). ((3x)/(4)-(1)/(2))-((x)/(4)-(3)/(2)) Multiply -1 by each term inside the parentheses. ((3x)/(4)-(1)/(2))-(x)/(4)+(3)/(2) Remove the parentheses around the expression (3x)/(4)-(1)/(2). (3x)/(4)-(1)/(2)-(x)/(4)+(3)/(2) Combine the numerators of all expressions that have common denominators. (3x-x)/(4)-(1)/(2)+(3)/(2) Combine all like terms in the numerator. (2x)/(4)-(1)/(2)+(3)/(2) Reduce the expression (2x)/(4) by removing a factor of 2 from the numerator and denominator. (x)/(2)-(1)/(2)+(3)/(2) Combine the numerators of all expressions that have common denominators. (x-1)/(2)+(3)/(2) Combine the numerators of all expressions that have common denominators. ((x-1)+3)/(2) Combine all like terms in the numerator. (x+2)/(2)

OpenStudy (vishweshshrimali5):

\[(\frac{3}{4} * x - \frac{1}{2}) - (\frac{1}{4} * x - \frac{3}{2}) = \frac{2}{4} * x + \frac{2}{2} = \frac{1}{2}*x + 1\]

OpenStudy (vishweshshrimali5):

\[= \frac{x}{2} + 1 = \frac{x+2}{2}\]